OpenStudy (anonymous):

-13 = 2b - b - 10 (please show work )

OpenStudy (anonymous):

-13+10=b -3=b

OpenStudy (anonymous):

how did u get it though ?

OpenStudy (anonymous):

-13=2b-b-10 -13+10=2b-b -3=b

OpenStudy (anonymous):

thanks!